OpenStudy (anonymous):

What is the slope of the line given by the equation below? If necessary, use the slash mark ( / ) for a fraction bar. Y=6/5x

OpenStudy (mathmale):

Best to write this equation as (6/5) x, to emphasize that 6/5 is your slope. You could compare y = (6/5) x to y = m x + b and see by inspection that m=the slope=6/5 and b=the y intercept)=0.